# Abelian surfaces over finite fields containing no curves of genus $3$ or less

### v1

@@ Abstract.
We characterise abelian surfaces defined over finite fields containing no curves of genus less than or equal to $3$. Firstly, we complete and expand the characterisation of isogeny classes of abelian surfaces with no curves of genus up to $2$ initiated by the first author et al. in previous work. Secondly, we show that, for simple abelian surfaces, containing a curve of genus $3$ is equivalent to admitting a polarisation of degree $4$. Thanks to this result, we can use existing algorithms to check which isomorphism classes in the isogeny classes containing no genus $2$ curves have a polarisation of degree $4$. Thirdly, we characterise isogeny classes of abelian surfaces with no curves of genus $\leq 2$, containing no abelian surface with a polarisation of degree $4$. Finally, we describe absolutely irreducible genus $3$ curves lying on abelian surfaces containing no curves of genus less than or equal to $2$, and show that their number of rational points is far from the Serre–Weil bound.
@@ Introduction
Studying the minimal genus of algebraic curves lying on an abelian variety is a classical question [1, 8, 4]. In this paper we focus our attention on abelian surfaces. Over an algebraically closed field, every abelian surface is isogenous to a principally polarised one [20, Cor. 1, p. 234], that is, to the Jacobian of a genus 2 curve or to the product of two elliptic curves. A well-known consequence of this fact is that every abelian surface over an algebraically closed field contains a (possibly singular) absolutely irreducible curve of geometric genus $2$ or $1$; see also Lemma 1.1 below. On another note, one can also study when an abelian surface contains curves of a fixed genus. Over the complex numbers, this question has been treated in [5] for genus $3$ curves and in [7] for hyperelliptic curves of genus $4$. Over non-algebraically closed fields, the situation is more complicated. For example, it is no longer true that any abelian surface is isogenous to a principally polarised one.
From now on, we direct our attention to the case of finite fields. Isogeny classes in which no abelian surface admits a principal polarisation are classified in [14]. Furthermore, even if an isogeny class is principally polarisable, it might not contain a Jacobian surface nor the product of two elliptic curves [15]. Indeed, by a theorem of Weil [15, Thm. 1.3], we have one more option among principally polarisable abelian surfaces, that is, Weil restrictions of elliptic curves defined over a quadratic extension of the finite field. To sum up, studying the minimal genus of algebraic curves lying over an abelian surface defined over a finite field is a hard question, and it is in fact open in its full generality. Besides the intrinsic theoretical interest, this question was raised in the context of algebraic geometry coding theory. Indeed, the first author and her co-authors showed in [2] that so called algebraic geometry codes constructed from abelian surfaces defined over a finite field and containing no curves of small genus respect a better bound on their minimum distance.
@@
Let $\mathbb{F}_{q}$ be a finite field of cardinality $q$ the power of a prime $p$. We recall that an isogeny class of abelian surfaces over $\mathbb{F}_{q}$ corresponds via the Honda–Tate theory to a polynomial of the form $f(t)=t^{4}+at^{3}+bt^{2}+aqt+q^{2}$, called Weil polynomial.
Previous work by the first author and coauthors showed that abelian surfaces defined over $\mathbb{F}_{q}$ which are not isogenous to a principally polarised abelian surface or which are isogenous to Weil restrictions of some elliptic curves defined over $\mathbb{F}_{q^{2}}$ do not contain absolutely irreducible curves of arithmetic genus up to $2$ [2, Prop. 4.2 and 4.3]. We denote the set of Weil polynomials corresponding to the former by $\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}$ while the latter splits into $\{(t^{2}-2)^{2},(t^{2}-3)^{2}\}\sqcup\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$. We expand the previous result in Theorem 1.2, by showing that these abelian surfaces do not contain absolutely irreducible curves of geometric genus up to $2$, and that this is an equivalence. In Remark 1.3 we show that restricting to absolutely irreducible curves of arithmetic genus up to $2$ does not provide an equivalence.
The main goal of the present paper is to determine when an abelian surface $A$ does not contain curves of genus $\leq 3$. Our first main result is the following.
Main Theorem 1 (Theorem 2.3).
@@
As the notation suggests, the Weil polynomials of the aforementioned isogeny classes are all irreducible except for $(t^{2}-2)^{2}$ and $(t^{2}-3)^{2}$. The isogeny classes determined by $(t^{2}-2)^{2}$ and $(t^{2}-3)^{2}$ are dealt in Proposition 2.1. So we now assume that $f(t)$ is irreducible, that is $f(t)\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}\sqcup\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$. When the corresponding isogeny class $\mathcal{C}$ is ordinary, the third author designed an algorithm in [18] to compute the isomorphism classes of abelian surface in $\mathcal{C}$ admitting a polarisation of degree $4$. See Section 6 for an overview and examples.
Our second main result gives a complete classification of the isogeny classes determined by $f(t)\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}\sqcup\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$ that contain an abelian surface with a polarisation of degree $4$. The proof of this classification builds on Howe’s seminal work [12, 13] on kernels of polarisations of abelian varieties over finite fields. An expanded statement appears later in the paper as Theorem 5.7.
@@
    Assume $\mathcal{C}$ is ordinary. Then (italic-$\bigstar$2) holds if and only if there is no $A$ in $\mathcal{C}$ with maximal $\mathbb{F}_{q}$-endomorphism ring admitting a polarisation of degree $4$.
@@
    Assume that $f\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}$. Then (italic-$\bigstar$2) holds if and only if $2$ is inert in $K^{+}$.
@@
    Assume that $f(t)\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$, so that $f(t)=t^{4}+bt^{2}+q^{2}$. Then (italic-$\bigstar$2) is equivalent to
@@
In Section 1 we recall and complete the characterisation given in [2] of isogeny classes of abelian surfaces containing no absolutely irreducible curves of genus smaller than or equal to $2$, and describe some properties of their Weil polynomials. In Section 2, we start the characterisation of abelian surfaces containing no curves of genus $3$ among those containing no curves of genus up to $2$. In particular, in Theorem 2.3, we prove the key equivalence between containing a curve of arithmetic genus $3$ and having a polarisation of degree $4$. In Section 3 we collect technical results on the factorisation of $2$ in the extension $K/\mathbb{Q}$, that we shall use throughout the paper. Section 4 is devoted to results on kernels of polarisations of abelian varieties, based on the work of Howe [12, 13]. In Section 5 we prove our main Theorem 5.7, giving necessary and sufficient conditions on an isogeny class to not contain surfaces admitting a polarisation of degree $4$, hence containing no curves of arithmetic genus up to $3$. Section 6 describes the algorithm proposed by the third author [18] to compute the isomorphism classes of abelian surface in an isogeny class admitting a polarisation of degree $4$, and provides examples for our main Theorem 5.7. Finally, in Section 7, we characterise curves of genus $3$ lying on abelian surfaces containing no curves of genus less than $2$, and give bounds for their number of rational points, showing that those genus $3$ curves are far from being maximal.
@@ 1. Abelian surfaces containing no curves of genus $\leq 2$
Let $A$ be a $g$-dimensional abelian variety defined over the finite field $\mathbb{F}_{q}$ of characteristic $p$. The characteristic polynomial $f(t)$ of its Frobenius endomorphism acting on the $\ell$-Tate module (for any prime $\ell\neq p$) is a monic polynomial of degree $2g$ with integer coefficients. All the complex roots of $f(t)$ have absolute value $\sqrt{q}$. We call such a polynomial a Weil polynomial. Honda and Tate showed in [25, 11, 26] that $f(t)$ completely determines the isogeny class $\mathcal{C}$ of $A$. Recall that an abelian variety $A$ over $\mathbb{F}_{q}$ of dimension $g$ is called ordinary if the coefficient of $t^{g}$ in $f(t)$ is coprime with $q$. In particular, being ordinary is a property of the isogeny class $\mathcal{C}$ of $A$. We will say that $\mathcal{C}$ and $f(t)$ are ordinary if $A$ is so.
@@
By a curve $C$ we mean a possibly singular one-dimensional $\mathbb{F}_{q}$-irreducible projective variety always defined over $\mathbb{F}_{q}$. We say that a curve is absolutely irreducible if it is irreducible over the algebraic closure $\overline{\mathbb{F}}_{q}$ of $\mathbb{F}_{q}$. Nevertheless, later in the paper, when we want to emphasise that a curve $C$ is not necessarily absolutely irreducible, we will say that $C$ is $\mathbb{F}_{q}$-irreducible. We say that a curve $C$ lies on an abelian surface $A$ or that $A$ contains $C$, if $C$ is a closed subvariety of $A$, or, equivalently, if there is an embedding of $C$ in $A$. We recall here some well-known definitions and results, following [23, Ch. 4, Sec. 2]. The arithmetic genus $p_{a}$ of a curve is defined as $1-\chi(C)$, $\chi(C)$ being the Euler-Poincaré characteristic of $C$. The geometric genus $g$ of $C$ is by definition the genus of its normalisation $\tilde{C}$. The geometric genus of a curve is always smaller than or equal to its arithmetic genus, and equality holds if and only if the curve is smooth. A divisor of $A$ is a formal sum, with integer coefficients, of $\mathbb{F}_{q}$-irreducible curves lying on $A$. It is called effective if all its coefficients are non-negative. By the adjunction formula [23, Ch. 4, Sec. 2, Prop. 5], since the class of canonical divisor of an abelian surface is trivial, for any $\mathbb{F}_{q}$-irreducible curve $C$ of arithmetic genus $p_{a}$ on $A$ we have $C^{2}=2p_{a}-2$.
The goal of this section is to prove some characterisations and properties of Weil polynomials of abelian surfaces which do not contain absolutely irreducible curves of geometric genus 00, $1$ or $2$. We start with the following well-known lemma.
@@
Let $f:A\to B$ be an isogeny of abelian varieties defined over a field $k$. Let $C$ be a curve of geometric genus $g$ defined over $k$, lying on $A$. Then, $f(C)$ is a curve of geometric genus $g$ on $B$. Moreover, if $C$ is absolutely irreducible then $f(C)$ is absolutely irreducible as well.
@@
It follows from the fact that $f(C)$ is birationally equivalent to $C$. ∎
Some of the implications of the next theorem are proved in [2]. However, as we will discuss in Remark 1.3, there were some gaps that we fill in.
@@
  3. (3)
    exactly one of the following statements hold:
@@
      $A$ is isogenous to a Weil restriction of an elliptic curve defined over the quadratic extension of $\mathbb{F}_{q}$ (hence $a=0$), which is equivalent to having $A\otimes_{\mathbb{F}_{q}}\mathbb{F}_{q^{2}}$ not simple, and one of the following conditions hold:
@@
      • $p\equiv 11\bmod 12$ or $p=3$, $q$ is a square and $b=-q$;
@@
[…]
@@
[…]
Finally, let us remark that the implication $(c)\Rightarrow(b)$ of [2, Lemma 4.1] still holds true, but a new proof is necessary, that we offer here. Let us suppose that $A$ contains a curve of arithmetic genus $p_{a}=1$. Then the geometric genus is necessarily $1$ as well, since $g$ cannot be 00 and $g\leq p_{a}$, thus the curve is smooth, a contradiction. Suppose now that $A$ contains a curve of arithmetic genus $p_{a}=2$. If the geometric genus is $1$ then we know by [9, Prop. 1] that the curve is smooth and has a structure of an elliptic curve, which leads to a contradiction. If the geometric genus is $2$ then the curve is smooth of genus $2$, leading again to a contradiction.

### v3

@@ Abstract.
We study abelian surfaces defined over finite fields which do not contain any possibly singular curve of genus less than or equal to $3$. Firstly, we complete and expand the characterisation of isogeny classes of abelian surfaces with no curves of genus up to $2$ initiated by the first author et al. in previous work. Secondly, we show that, for simple abelian surfaces, containing a curve of genus $3$ is equivalent to admitting a polarisation of degree $4$. Thanks to this result, we can use existing algorithms to check which isomorphism classes in the isogeny classes containing no genus $2$ curves have a polarisation of degree $4$. Thirdly, we characterise isogeny classes of abelian surfaces with no curves of genus $\leq 2$, containing no abelian surface with a polarisation of degree $4$. Finally, we describe the absolutely irreducible genus $3$ curves lying on abelian surfaces containing no curves of genus less than or equal to $2$, and show that their number of rational points is far from the Serre–Weil bound.
@@ Introduction
Studying the minimal genus of algebraic curves lying on an abelian variety is a classical question [1, 8, 4]. In this paper, we focus our attention on abelian surfaces. Over an algebraically closed field, every abelian surface is isogenous to a principally polarised one [22, Cor. 1, p. 234], that is, to the Jacobian of a smooth genus 2 curve or to the product of two elliptic curves. A well-known consequence of this fact is that every abelian surface over an algebraically closed field contains a (possibly singular) irreducible curve of geometric genus $2$ or $1$; see also Lemma 1.1 below. On another note, one can also study when an abelian surface contains curves of a fixed genus. Over the complex numbers, this question has been treated in [5] for genus $3$ curves and in [7] for hyperelliptic curves of genus $4$. Over non-algebraically closed fields, the situation is more complicated. For example, it is no longer true that any abelian surface is isogenous to a principally polarised one.
From now on, we direct our attention to the case of finite fields. Isogeny classes in which no abelian surface admits a principal polarisation are classified in [15]. Furthermore, even if an isogeny class is principally polarisable, it might not contain a Jacobian surface nor the product of two elliptic curves [16]. Indeed, by a theorem of Weil [16, Thm. 1.3], we have one more option among principally polarisable abelian surfaces, that is, Weil restrictions of elliptic curves defined over a quadratic extension of the finite field. To sum up, studying the minimal genus of algebraic curves lying over an abelian surface defined over a finite field is a hard question, and it is in fact open in its full generality.
Besides the intrinsic theoretical interest, this question was raised in the context of coding theory. The first author and her co-authors studied in [2] the so-called algebraic geometry codes constructed from a rational divisor on an abelian surface defined over a finite field $\mathbb{F}_{q}$. They showed that the minimum distance of such codes respects, under some hypotheses, a lower bound that is somewhat proportional to the minimum genus of the $\mathbb{F}_{q}$-irreducible curves defined over $\mathbb{F}_{q}$ lying on the surface [2, Thm. 3.3 and Rmk. 3.6]. Since codes with large minimum distance are desirable, this led to the question on how to find abelian surfaces containing no curves of low genus.
@@
In the following, by curve we mean an algebraic projective one-dimensional variety defined over a field. Note that for us a curve doesn’t need to be smooth or irreducible. In the rest of the paper, unless otherwise specified, we will assume the base field to be $\mathbb{F}_{q}$, a finite field with $q$ elements of characteristic $p$, and the curve to be $\mathbb{F}_{q}$-irreducible. The definition of arithmetic and geometric genus of such a curve is recalled at the beginning of Section 1.
An isogeny class of abelian surfaces over $\mathbb{F}_{q}$ corresponds via the Honda–Tate theory to a polynomial of the form $f(t)=t^{4}+at^{3}+bt^{2}+aqt+q^{2}$ in $\mathbb{Z}[x]$, called Weil polynomial. Previous work by the first author and coauthors showed that abelian surfaces defined over $\mathbb{F}_{q}$ which do not contain absolutely irreducible curves of arithmetic genus up to $2$ are either not isogenous to a principally polarised abelian surface or are isogenous to Weil restrictions of some elliptic curves defined over $\mathbb{F}_{q^{2}}$; see [2, Prop. 4.2 and 4.3]. We denote the set of Weil polynomials corresponding to the former isogeny classes by $\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}$ while the latter classes splits into $\{(t^{2}-2)^{2},(t^{2}-3)^{2}\}\sqcup\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$; see Definition 1.9 for the precise definition of these classes. As the notation suggests, the Weil polynomials of the aforementioned isogeny classes are all irreducible except for $(t^{2}-2)^{2}$ and $(t^{2}-3)^{2}$. These results are expanded in Main Theorem 1.
Main Theorem 1.
Let $f(t)$ be a Weil polynomial corresponding to an isogeny class $\mathcal{C}$ of simple abelian surfaces over $\mathbb{F}_{q}$. The following statement holds:
  1. (i)
    No $A$ in $\mathcal{C}$ contains an absolutely irreducible curve of geometric genus $\leq 2$ if and only if $f(t)$ is in $\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}\sqcup\{(t^{2}-2)^{2},(t^{2}-3)^{2}\}\sqcup\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$.
  2. (ii)
    No $A$ in $\mathcal{C}$ contains an $\mathbb{F}_{q}$-irreducible curve of arithmetic genus $\leq 2$ if and only if $f(t)$ is in $\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}$.
  3. (iii)
    No $A$ in $\mathcal{C}$ contains an absolutely irreducible curve with geometric genus $\leq 2$ but there exists a $B$ in $\mathcal{C}$ which contains an $\mathbb{F}_{q}$-irreducible curve with arithmetic genus $2$ if and only if $f(t)$ is in $\{(t^{2}-2)^{2},(t^{2}-3)^{2}\}\sqcup\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$.
An expanded version of Part i is in Theorem 1.4, where we also give a characterisation of the Weil polynomials in terms of their coefficients. In Remark 1.5 we show that restricting the statement of Part i to absolutely irreducible curves of arithmetic genus up to $2$ does not provide an equivalence. Parts ii and iii are proven as Corollary 1.10.
The rest of the paper aims at extending Main Theorem 1 to curves of genus $\leq 3$. The main tool for this purpose will be our second main result, which can be found later in the text as Theorem 2.2.
Main Theorem 2.
@@
Firstly, we treat the isogeny classes $\mathcal{C}$ and $\mathcal{C}^{\prime}$ defined by the reducible Weil polynomials $(t^{2}-2)^{2}$ and $(t^{2}-3)^{2}$, respectively. As we know by Main Theorem 1.iii, both $\mathcal{C}$ and $\mathcal{C}^{\prime}$ contain at least one abelian surface with an $\mathbb{F}_{q}$-irreducible – but not absolutely irreducible – curve of arithmetic genus $2$ lying on it. So our focus is on absolutely irreducible curves of arithmetic genus $3$. In Proposition 2.4, we show that no abelian surface in $\mathcal{C}$ contains an absolutely irreducible curve of geometric or arithmetic genus $3$, while there exists an abelian surface in $\mathcal{C}^{\prime}$ containing an absolutely irreducible smooth curve of genus $3$.
Secondly, we consider the irreducible Weil polynomials $f(t)\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}\sqcup\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$. When the corresponding isogeny class $\mathcal{C}$ is ordinary, the third author designed an algorithm in [19] to compute the isomorphism classes of abelian surface in $\mathcal{C}$ admitting a polarisation of degree $4$, and hence containing an $\mathbb{F}_{q}$-irreducible curve of arithmetic genus $3$. See Section 6 for an overview and examples. By applying Main Theorem 2 to our isogeny classes, we easily obtain the following statements about curves of arithmetic genus $\leq 3$ (and not just $=3$). See Corollary 2.5 for the proof.
Corollary.
Let $f(t)$ be the Weil polynomial of a simple isogeny class $\mathcal{C}$ of abelian surfaces over $\mathbb{F}_{q}$.
  1. (i)
    Assume that $f(t)\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}\sqcup\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$ and let $A$ be an abelian surface in $\mathcal{C}$. If $A$ contains an absolutely irreducible curve of arithmetic genus $\leq 3$ then $A$ admits a polarisation of degree $4$.
  2. (ii)
    No abelian surface in $\mathcal{C}$ contains an $\mathbb{F}_{q}$-irreducible curve of arithmetic genus $\leq 3$ if and only if $f(t)$ is in $\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}$ and no abelian surface in $\mathcal{C}$ admits a polarisation of degree $4$.
Finally, in Main Theorem 3 below, we give a complete classification of the isogeny classes determined by $f(t)\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}\sqcup\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$ that contain an abelian surface with a polarisation of degree $4$. The proof of this classification builds on Howe’s seminal work [13, 14] on kernels of polarisations of abelian varieties over finite fields. An expanded version of Main Theorem 3 appears later in the paper as Theorem 5.7.
@@
    Assume $\mathcal{C}$ is ordinary. Then (italic-$\bigstar$3) holds if and only if there is no $A$ in $\mathcal{C}$ with maximal $\mathbb{F}_{q}$-endomorphism ring admitting a polarisation of degree $4$.
@@
    Assume that $f(t)\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}$. Then (italic-$\bigstar$3) holds if and only if $2$ is inert in $K^{+}$.
@@
    Assume that $f(t)\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$, so that $f(t)=t^{4}+bt^{2}+q^{2}$. Then (italic-$\bigstar$3) is equivalent to
@@
In Section 1, we prove Main Theorem 1, expanding on [2] and characterising the isogeny classes of abelian surfaces containing no absolutely irreducible curves of geometric genus smaller than or equal to $2$. In Section 2, we start the characterisation of abelian surfaces containing no curves of genus $3$ among those containing no curves of genus up to $2$. In particular, in Theorem 2.2, we prove the key equivalence between containing a curve of arithmetic genus $3$ and having a polarisation of degree $4$ stated above as Main Theorem 2. We derive also some consequences stated above in the Corollary. In Section 3 we collect technical results on the factorisation of $2$ in the extension $K/\mathbb{Q}$, that we shall use throughout the paper. Section 4 is devoted to results on kernels of polarisations of abelian varieties, based on the work of Howe [13, 14]. In Section 5, we prove an expanded version of Main Theorem 3, namely Theorem 5.7, in which we give necessary and sufficient conditions on an isogeny class to not contain surfaces admitting a polarisation of degree $4$. Section 6 describes the algorithm proposed by the third author in [19] to compute the isomorphism classes of abelian surface in an isogeny class admitting a polarisation of degree $4$, and provides examples for our Theorem 5.7. Finally, in Section 7, we characterise absolutely irreducible smooth curves of genus $3$ lying on abelian surfaces containing no curves of genus less than $2$, and give bounds for their number of rational points, showing that such genus $3$ curves are far from being maximal.
@@ 1. Abelian surfaces containing no curves of genus $\leq 2$
Let $A$ be a $g$-dimensional abelian variety defined over a finite field $\mathbb{F}_{q}$ of characteristic $p$. The characteristic polynomial $f(t)$ of its Frobenius endomorphism acting on the $\ell$-Tate module (for any prime $\ell\neq p$) is a monic polynomial of degree $2g$ with integer coefficients. All the complex roots of $f(t)$ have absolute value $\sqrt{q}$. We call such a polynomial a Weil polynomial. Honda and Tate showed in [27, 11, 28] that $f(t)$ completely determines the isogeny class $\mathcal{C}$ of $A$. Recall that an abelian variety $A$ over $\mathbb{F}_{q}$ of dimension $g$ is called ordinary if the coefficient of $t^{g}$ in $f(t)$ is coprime with $q$. In particular, being ordinary is a property of the isogeny class $\mathcal{C}$ of $A$. We will write that $\mathcal{C}$ and $f(t)$ are ordinary if $A$ is so.
@@
By a curve we mean a possibly singular one-dimensional projective variety defined over $\mathbb{F}_{q}$. We will only consider $\mathbb{F}_{q}$-irreducible curves in this paper. In some cases, we will need our curves to be absolutely irreducible, that is, irreducible over the algebraic closure $\overline{\mathbb{F}}_{q}$ of $\mathbb{F}_{q}$. We say that a curve $C$ defined over $\mathbb{F}_{q}$ lies on an abelian surface $A$ or that $A$ contains $C$, if $C$ is a closed subvariety of $A$, that is, there is an $\mathbb{F}_{q}$-embedding of $C$ in $A$. A divisor of $A$ defined over $\mathbb{F}_{q}$, or a rational divisor over $\mathbb{F}_{q}$, is a formal sum, with integer coefficients, of $\mathbb{F}_{q}$-irreducible curves defined over $\mathbb{F}_{q}$ and lying on $A$ [^1] A divisor is often called rational over $\mathbb{F}_{q}$ when it is invariant under the action of $\mathrm{Gal}(\overline{\mathbb{F}}_{q}/\mathbb{F}_{q})$. This implies in particular that the divisor can be written as a formal sum of $\mathbb{F}_{q}$-irreducible curves defined over $\mathbb{F}_{q}$ (see [9, §3.1]). In the present paper we shall only use this characterization, hence we take it as a definition.. It is called effective if all its coefficients are non-negative and ample if a multiple of it is very ample, that is, defines an embedding of the surface in a projective space.
The arithmetic genus $p_{a}(C)$ of an $\mathbb{F}_{q}$-irreducible curve $C$ is defined as $1-\chi(C)$, $\chi(C)$ being the Euler-Poincaré characteristic of $C$ [25, Ch. IV, Sec. 2]. Consider the absolutely irreducible components $C_{i}$ of $C$, and their normalisations $\tilde{C_{i}}$. We define the geometric genus $g(C)$ of $C$ as $g(C)=p_{a}(\tilde{C_{1}})+\ldots+p_{a}(\tilde{C_{n}})$. We have that $g(C)\leq p_{a}(C)$ with an equality if and only if $C$ is smooth. If $C$ is an $\mathbb{F}_{q}$-irreducible curve lying on an abelian surface $A$ then, since the class of the canonical divisor of $A$ is trivial, we have that $C^{2}=2p_{a}(C)-2$ by the adjunction formula [25, Ch. IV, Sec. 2, Prop. 5].
The goal of this section is to prove some characterisations and properties of Weil polynomials of abelian surfaces which do not contain absolutely irreducible curves of geometric genus 00, $1$ or $2$. We start with the following useful lemma.
@@
Let $f\colon A\to B$ be an isogeny of abelian varieties defined over $\mathbb{F}_{q}$. Let $C$ be an $\mathbb{F}_{q}$-irreducible curve of geometric genus $g$ defined over $\mathbb{F}_{q}$, lying on $A$. Then, $f(C)$ is an $\mathbb{F}_{q}$-irreducible curve of geometric genus $\leq g$ on $B$. Moreover, if $C$ is absolutely irreducible then $f(C)$ is absolutely irreducible as well.
@@
[…]
For the final statement, assume that $C$ is absolutely irreducible and let $\bar{C}$ and $\bar{f}$ be the extension of $C$ and $f$, respectively, to an algebraic closure $\bar{\mathbb{F}}_{q}$ of $\mathbb{F}_{q}$. Then $\bar{C}$ is irreducible, and since $\bar{f}$ is continuous, we get that $\bar{f}(\bar{C})$ is irreducible as well. Since $\bar{f}(\bar{C})$ is the extension of $D$ to $\bar{\mathbb{F}}_{q}$, we conclude that $D$ is absolutely irreducible. ∎
The statement and proof of Lemma 1.1 carry over verbatim if we replace $\mathbb{F}_{q}$ by an arbitrary perfect field.
The following results relate the existence of a principal polarisation on an abelian surface $A$ and of curves of genus $2$ on $A$.
Lemma 1.2.
Let $C$ and $D$ be two $\mathbb{F}_{q}$-irreducible curves lying on an abelian surface $A$ such that $C.D=0$. Then, $C$ and $D$ are both elliptic curves lying on $A$.
Proof.
This is part of [5, Lemma 1.1] whose proof adapts to finite fields without any change. ∎
Proposition 1.3.
Let $A$ be a simple abelian surface over $\mathbb{F}_{q}$. Then the following statements are equivalent:
  1. (i)
    $A$ has a principal polarisation;
  2. (ii)
    $A$ contains an $\mathbb{F}_{q}$-irreducible curve of arithmetic genus $2$.
Proof.
Assume that $A$ has a principal polarisation. Since $\mathbb{F}_{q}$ is finite, there exists an ample invertible sheaf $\mathcal{L}\in\mathrm{Pic}(A)$ defined over $\mathbb{F}_{q}$ such that the principal polarisation is given by $\lambda_{\mathcal{L}}$ [20, Remark 13.2]. By [22, p. 150], if we write $\mathcal{L}=\mathcal{L}(D)$ for some effective rational divisor $D$ then
(1) $\deg\lambda_{\mathcal{L}}=\chi(D)^{2}=(D^{2}/2)^{2}.$
Since $\deg\lambda_{\mathcal{L}}=1$, it follows that $D^{2}=2$. Hence, if $D=E+F$ were reducible then, from the formula
(2) $D^{2}=E^{2}+F^{2}+2\cdot E.F,$
we would get that $D$ is the sum of two elliptic curves. Since $A$ is simple, it does not contain any elliptic curve. Hence, $D$ is an $\mathbb{F}_{q}$-irreducible curve of arithmetic genus $2$.
Conversely, assume that $A$ contains an $\mathbb{F}_{q}$-irreducible curve $C$ of arithmetic genus $2$. By the adjunction formula, we get $C^{2}=2$. Lemma 1.2 implies that $C$ intersects any other $\mathbb{F}_{q}$-irreducible curve on $A$ positively. Hence, the Nakai–Moishezon criterion [10, Sec. 5, Thm. 1.10] says that the divisor defined by $C$ is ample. So, the induced isogeny $A\to A^{\vee}$ is a polarisation. Therefore, $C$ defines a principal polarisation on $A$ by Equation (1). ∎
Part of the statement of the next theorem is proved in [2]. However, as we will discuss in Remark 1.5, there were some gaps that we fill in.
@@
  3. (iii)
    exactly one of the following statements holds:
@@
      $A$ is isogenous to a Weil restriction of an elliptic curve defined over the quadratic extension of $\mathbb{F}_{q}$ (hence $a=0$) and one of the following conditions holds:
@@
      • $p\equiv 11\bmod 12$, $q$ is a square and $b=-q$;
      • $p=3$, $q$ is a square and $b=-q$;
@@
We start by proving (i)$\Rightarrow$(ii). By assumption, $A$ does not contain an elliptic curve, hence it is simple. Assume that $A$ is isogenous to the Jacobian of a smooth absolutely irreducible genus $2$ curve $C$. Since $C$ is canonically embedded into its Jacobian, Lemma 1.1 would imply that $A$ contains an absolutely irreducible curve of geometric genus $\leq 2$, in contradiction with (i).
Now we show that (ii)$\Rightarrow$(i). First note that $A$ cannot contain an absolutely irreducible curve $D$ of geometric genus 00. Indeed, the normalisation $\tilde{D}$ of $D$ is birationally equivalent to $\mathbb{P}^{1}$ and, by [20, Cor. 3.8], the only rational maps from $\mathbb{P}^{1}$ to $A$ are the constant maps. The abelian variety $A$ cannot contain an absolutely irreducible curve of geometric genus $1$ as well. Indeed, the normalisation of such a curve would be an elliptic curve (see for example [9, Prop. 1]) and hence an isogeny factor of $A$. Similarly, $A$ cannot contain an absolutely irreducible curve $D$ of geometric genus $2$ because the induced map $\tilde{D}\to A$ would give a morphism $\mathop{\rm Jac}\nolimits(\tilde{D})\to A$, which is an isogeny since $A$ is simple.
We now focus on (ii)$\Rightarrow$(iii). We distinguish two cases. Assume first that $A$ is not isogenous to a principally polarised abelian surface. This corresponds to (iii)(iii).(a). The equivalence in terms of the coefficients of the Weil polynomial follows from [15, Thm. 1]. Assume now that $A$ is isogenous to a principally polarised abelian surface. By a classification theorem due to Weil (see for instance [16, Thm. 1.3]), a principally polarised abelian surface defined over $\mathbb{F}_{q}$ is exactly one of the following: a product of two elliptic curves defined over $\mathbb{F}_{q}$ with the product polarisation; the Jacobian of an absolutely irreducible genus $2$ curve defined over $\mathbb{F}_{q}$ with the canonical polarisation; the Weil restriction of an elliptic curve defined over the quadratic extension of $\mathbb{F}_{q}$ with the induced polarisation. Since we are assuming (ii), we exclude the first two cases. So, $A$ is the Weil restriction of an elliptic curve defined over the quadratic extension of $\mathbb{F}_{q}$. Finally, the characterisation of the coefficients of the Weil polynomial follows from [2, Prop. 4.3] and the beginning of its proof. Hence, we are in case (iii)(iii).(b).
We conclude by showing that (iii)$\Rightarrow$(ii). For $A$ satisfying (iii)(iii).(a) the implication is clear. If $A$ satisfies (iii)(iii).(b) the result follows from [16, Thm. 1.2-(2), Table 1.2]. ∎
Remark 1.5.
In [2, Lemma 4.1], it is stated that for an abelian surface $A$ defined over $\mathbb{F}_{q}$ we have that the following statements are equivalent:
  1. (I)
    $A$ is simple and not isogenous to a Jacobian surface, that is, the Jacobian of an absolutely irreducible smooth genus $2$ curve;
  2. (II)
    $A$ does not contain absolutely irreducible curves of arithmetic genus 00, $1$ or $2$;
  3. (III)
    $A$ does not contain absolutely irreducible smooth curves of genus 00, $1$ or $2$.
@@
We highlight that there is a crucial difference between considering the arithmetic genus and the geometric genus in our characterisation. On the one hand, Lemma 1.1 implies that if an abelian surface $A$ doesn’t contain absolutely irreducible curves of geometric genus $\leq 3$ then every abelian variety isogenous to $A$ doesn’t contain absolutely irreducible curves of geometric genus $\leq 3$. On the other hand, this is not the case for absolutely irreducible curves of arithmetic genus $\leq 2$ as we have seen in Remark 1.5, or for absolutely irreducible curves of arithmetic genus $\leq 3$ as we will see using Theorem 2.2.